OpenStudy (anonymous):

Help me Find the horizontal asymptote of f(x)=ln(x^2sin^2(x))-ln(x^2+4)

OpenStudy (alekos):

it might pay to use desmos to graph this function

OpenStudy (tkhunny):

Have you considered properties of logarithms? \(x^{2} + 4 \ne 0\) \(f(x) = ln\left(\sin^{2}(x)\dfrac{x^{2}}{x^{2}+4}\right)\) Seems a bit more obvious in that form.

OpenStudy (anonymous):

ln(a)-ln(b) = ln(a/b)

OpenStudy (anonymous):

Then how I can continue to get a solution